# Happy family of stable marriages

### 02 Apr 2026 v1

@@ Abstract
This paper introduces a natural definition for the volume of the unit ball in $n$-dimensional normed spaces $\mathbb{R}^{n}$. This definition preserves the Euclidean relation $P(B)/V(B)=n$ between the perimiter and the volume of the unit ball $B$ in $R^{n}$. We show that this volume definition is invariant under origin-preserving affine transformations and polar duality. For $n=2$, we derive an explicit integral formula for the self-perimeter of the unit ball, extend it to non-centrally symmetric sets;. The construction is extended to $\mathbb{R}^{n}$ via a recursive integration over the boundary, utilizing (n1)(n-1)-dimensional volumes of planar intersections. Finally, we pose and discuss an Alexandrov-type problem for the associated surface measure, providing perturbative solutions in the 2D case. In particular we prove that, generically, any perturbation of the surface measure of the Euclidean 2-D disk yields a 4-fold symmetric convex set in the leading order.
@@
Technion, Israel Inst. of Thechnology, Haifa,Israel
@@ 1 Introduction
The concept of volume and surface area in finite-dimensional normed spaces (Minkowski spaces) lacks a single, universally accepted definition, unlike in standard Euclidean geometry. Over the years, various approaches have been proposed, each tailored to preserve specific geometric, analytic, or integral properties of the space [1, 5].
@@
Among the most prominent classical frameworks are the Busemann and the Holmes-Thompson definitions. The Busemann definition [2] normalizes the measure such that the Minkowski unit ball has the same volume as the Euclidean unit ball. The corresponding surface area is then obtained by integrating the Euclidean surface element, normalized by the volume of the central cross-section of the unit ball parallel to the tangent plane. Alternatively, the Holmes-Thompson definition [3], originating from symplectic geometry, measures the volume of a set by multiplying its Lebesgue measure with the Lebesgue measure of the polar body in the dual space, normalized by the Euclidean unit ball volume. This approach inherently satisfies the Blaschke-Santaló inequalities [4] and is completely coordinate-independent.
However, these classical definitions do not inherently prioritize the preservation of the fundamental Euclidean relationship between the perimeter (or surface area) $P_{n}$ and the volume $V_{n}$ of the unit ball, namely $P_{n}=nV_{n}$.
In this paper, we attempt to find a natural definition of volume in normed spaces $\mathbb{R}^{n}$ based precisely on the preservation of this Euclidean ratio $P(B)/V(B)=n$ for the unit ball $B\subset\mathbb{R}^{n}$. To achieve this, we introduce an intrinsic, recursive approach, starting with $n=2$ and using the definition of self-perimeter
The concept of self-perimeter is a central object of study in asymmetric Minkowski geometry and the theory of convex distance functions. A fundamental extremal problem in this field involves bounding the minimum self-perimeter of convex bodies in dimension two. The historical starting point is Golab’s theorem [6], which established that for centrally symmetric convex bodies, the self-perimeter is bounded between six and eight (with the maximum attained by a parallelogram, the minimum by the perfect hexagon and the Euclidean $2\pi$ lies in between ).
@@
Extending this problem to the space of asymmetric convex bodies expands the upper bound. The precise result states that the minimum self-perimeter of any planar convex body (calculated with respect to the optimal interior reference point) is bounded between six and nine. The lower bound of six is attained if and only if the body is an affinely regular hexagon. The absolute maximum of nine is attained if and only if the convex body is a triangle, a property formally proven by Grünbaum [7]. These results are further supported by the strict convexity of the self-perimeter with respect to the location of the interior reference point [8].
For regular $k$-gons, as well as for the family of planar convex bodies possessing $2\pi/k$ rotational symmetry, the self-perimeter exhibits periodic behavior depending on the remainder of $k$ modulo four [9, 10]. The literature demonstrates that when $k\equiv 0\pmod{4}$, the regular polygon constitutes a maximum (upper bound) for the family of shapes with the same degree of symmetry. Conversely, when $k\equiv 2\pmod{4}$, the regular polygon serves as a minimum [11]. This modular partition dictates the nature of the convergence of the self-perimeter to the limit of $2\pi$ (the value for a circle).
Extension to higer dimensions
@@
In this paper we introduce a twist on Busemann definition [2] of a perimeter of a convex set $K$ containing the origin in its interior
@@
$\sigma_{K}^{n-1}(\nu_{x})\equiv\frac{\omega_{K}^{(n-1))}}{{\cal H}_{n-1}(K\cap\Sigma^{n-1}(\nu_{x}))}$
$\Sigma^{n-1}(\nu_{x})$ is the n1n-1 hyperplane perpendicular to the normal $\nu_{x}\in\partial K$ containing the origin and
$\omega_{n-1}=\frac{\pi^{n-1}/2}{\Gamma\left(\frac{n-1}{2}+1\right)}$
is the volume of the unit n1n-1 ball in Euclidean space.
@@
In our approach we replace $\omega_{n-1}$ in $\sigma_{K}(\nu_{x})$ by the a recursive definition of the volume of the n1n-1 unit ball defined by a center symmetric convex set $B\in\mathbb{R}^{n}$
$V(B^{(k)})=k^{-1}\int_{\partial B^{(k)}}\frac{V(B^{(k-1)}(\nu_{x}))d{\cal H}_{k-1}(x)}{{\cal H}_{k-1}(B^{(k-1)}(\nu_{x}))}\ \ k=2\ldots n$
where $B^{(k-1)}(\nu_{x}):=B^{(k)}\cap\Sigma^{k-1}(\nu_{x})$ and $V(B^{(1)}(\nu))\equiv 2$ for any $\nu\in\mathbb{S}^{1}$. Thus, the perimeter of a convex set $K$ in terms of the Minkowski norm $\|\cdot\|_{B}$ whose unit ball is $B$ is given by
$P(K)=\int_{\partial K}\frac{V(B^{n-1)}(\nu_{v}))}{{\cal H}_{n-1}(K\cap\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)\ ,$
namely we replace $\omega_{n-1}$ in (1) by $V(B^{(n-1)}(\nu_{v}))$ which is the self volume of the unit ball $B^{(n-1)}(\nu_{v})$ of the Minkowski norm on $\Sigma^{n-1}(\nu_{x})$ .
@@
Note that our definition coincides with Busemann definition for $n=2$ since $\omega_{1}=V(B^{(1)})=2$. However, it replaces Busemann’s ansatz of preservation of the Euclidean volume of the unit ball $\omega_{n}$ with the preservation of the Euclidean ratio $P(B)/V(B)=n$.
@@
The rational behind our definition is as follows: While Busemann definition of the perimeter of convex set scale the surface area by the Lebesgue measure of the parallel central cuts, we acknowledge the non isotropy induced by the norm by scaling it with respect to the projection of this norm on the subspaces. It is remarkable that while Busemann definition (1) does not show any connection with the Minkowski norm, it does give the self perimeter when applied to the unit ball of the same norm in dimension 2. This seems to be just a coincidence, due to the ansatz $V(B^{(1)})=\omega_{1}$.
@@
As we will show, the exact cancellation of the Jacobian in this formulation ensures that this definition of volume is strictly preserved under origin-preserving affine transformations, as well as under its polar image. In section 5 we extend this definition to non center symmetric convex bodies and prove that calculate explicitly the self perimeter of an $n$-simplex $\Delta^{n}$ and prove its convexity with respect to the chosen center. This implies the center-convexity of any convex body in $\mathbb{R}^{n}$ composed of Cartesian product of simplexes and 2D2-D convex sets. The center convexity of general convex body is conjectured,
@@
Beyond the foundational task of defining volume and surface area, this framework allows us to pose a novel inverse problem. In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively [4]. Within the context of our intrinsic definitions, we introduce an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set [1], [5].
Given that our perimeter and volume are defined recursively by integrating over the essential boundary-normalized by the self-volume of (n1)(n-1)-dimensional cross-sections-the natural inverse question arises: Given a positive measure $\mu$ on the Euclidean sphere $S^{n-1}$, does there exist a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\mu$?
In section 6 we explore this problem in the two-dimensional case, providing perturbative solutions and identifying necessary conditions for its solvability around the Euclidean disc. In particular we obtain that an $O(\epsilon)$ perturbation of the uniform measure on the unit Euclidean circle yields an $O(\epsilon^{1/2})$ perturbation of this circle, which depends only on the log- 4-fold symmetric part of the measure (c.f Theorem 6.1).
@@ 2 Fundamental definitions
Let $B\subset\mathbb{R}^{n}$ be a convex, center symmetric set (CCS) . The norm $\|\cdot\|_{B}$ is defined in the usual way as
@@
Let $K$ be a starlike set in $\mathbb{R}^{n}$ with respect the origin. For $\theta\in S^{(n-1)}$ set
$r_{K}(\theta)=\sup\left\{t>0;t\theta\in K\right\}\$
@@
If $K$ is center symmetric (CC) then $r_{K}(\theta)=r_{K}(-\theta)$.
Definition 2.2.
@@
3 Self perimeter in the 2-plane
If $n=2$ there is a natural notion of a perimeter $P_{B}$ of $B$ as measured with respect to this norm. Indeed, we can use the norm $\|\cdot\|_{B}$ as a natural definition of length of an arc.
In the case $n=2$ we can introduce an explicit formulation for the perimeter.
@@ 3 Self perimeter in the 2-plane
where $ds$ is the Euclidian length of the boundary of $B$, $\vec{\tau}(s)$ is the tangent vector to the boundary at $s\in\partial B$.
An explicit expression in terms of an integral on the Euclidean circle $\mathbb{S}^{1}$ is
@@
The self-perimeter of a convex $B\ni\{0\}$ in $\mathbb{R}^{2}$ is
$P(B)=\int_{0}^{2\pi}\frac{\sqrt{r^{2}_{B}(\phi)+(r^{{}^{\prime}}_{B}(\phi))^{2}}d\phi}{r_{B}\left(\phi+\arctan\left(\frac{r_{B}(\phi}{r_{B}^{{}^{\prime}}(\phi)})\right)\right)}$ (3)
where $r_{B}(\phi):=r_{B}(\cos(\phi),\sin(\phi))$
@@
The vector $\vec{r}=(r\cos(\phi),r\sin(\phi))$. Its tangent is $d\vec{r}/d\phi=(r^{{}^{\prime}}\cos(\phi)-r\sin(\phi),r^{{}^{\prime}}\sin(\phi)+r\cos(\phi)$. The slop of this direction is
@@
The self perimeter of CCS is alwasy between 4 and 6.
Corollary 3.1 (Golkab[6]).
If $r_{B}(\phi)$ is the radius function of a CCS in $\mathbb{R}^{2}$ then (3) takes values between 6 and 8.
Note that for $B\in CCS(\mathbb{R}^{2})$, $r_{B}$ is a $\pi$ periodic function.
The definition (3) can easily be extended to convex, non center symmetric sets. Let $K$ be such a set and $p\in int(B)$. We can define the radus
@@
If $B\ni p$ is convex but not center symmetric there is no upper limit to (3). Indeed, if $p\in B$ is very close to $\partial B$ then evidently the self perimeter can be very large.
The following results was established in [8]. For the sake of completeness we provide a proof below.
@@
Let $K\subset\mathbb{R}^{2}$ be a bounded convex set and $p\in Int(K)$. Then $p\rightarrow P(K-p$ is strongly convex in $Int(K$. In particular there exists a unique interior point $p\in\text{int}(K)$ that minimizes the self-perimeter $P(K-\{p\})$.
Proof.
The self-perimeter of $K$ with respect to an interior point $p$ is defined by the integral:
$P(p)=\int_{\partial K}\frac{1}{r_{K,p}(\vec{\tau}(s))}ds$
@@
Hence, $r_{K,p}(v)$ is a concave function of $p$.
@@
Since the self-perimeter $P(p)$ is an integral (a continuous sum) of strictly convex functions over the boundary $\partial K$ with respect to a positive length measure $ds$, the function $P(p)$ itself is strictly convex on $\text{int}(K)$.

### 09 Apr 2026 v2

@@ Abstract
This paper introduces a natural definition for the volume of the unit ball in $n$-dimensional normed spaces $\mathbb{R}^{n}$. This definition preserves the Euclidean relation $P(B)/V(B)=n$ between the perimeter and the volume of the unit ball $B$ in $\mathbb{R}^{n}$. We show that this volume definition is invariant under origin-preserving affine transformations. For $n=2$, we derive an explicit integral formula for the self-perimeter of the unit ball and extend it to non-centrally symmetric sets. The construction is extended to $\mathbb{R}^{n}$ via a recursive integration over the boundary, utilizing (n1)(n-1)-dimensional volumes of planar intersections, and we discuss its invariance under affine transformations. Finally, we pose and discuss an Alexandrov-type problem for the associated surface measure, providing perturbative solutions in the 2D case. In particular, we prove that, generically, any perturbation of the surface measure of the Euclidean 2D disk yields a 4-fold symmetric convex set in the leading order.
@@
Technion, Israel Inst. of Technology, Haifa, Israel
@@ 1 Introduction
The concept of volume and surface area in finite-dimensional normed spaces (Minkowski spaces) lacks a single, universally accepted definition, unlike in standard Euclidean geometry. Over the years, various approaches have been proposed, each tailored to preserve specific geometric, analytic, or integral properties of the space [1, 5]. Among the most prominent classical frameworks are the Busemann and the Holmes-Thompson definitions. The Busemann definition [2] normalizes the measure such that the Minkowski unit ball has the same volume as the Euclidean unit ball. The corresponding affine invariant surface area is then obtained by integrating the Euclidean surface element, normalized by the volume of the central cross-section of the unit ball parallel to the tangent plane. Alternatively, the Holmes-Thompson definition [3], originating from symplectic geometry, measures the volume of a set by multiplying its Lebesgue measure with the Lebesgue measure of the polar body in the dual space, normalized by the Euclidean unit ball volume. This approach inherently satisfies the Blaschke-Santaló inequalities [4] and is completely coordinate-independent. However, these classical definitions do not inherently prioritize the preservation of the fundamental Euclidean relationship between the perimeter (or surface area) $P_{n}$ and the volume $V_{n}$ of the unit ball, namely $P_{n}=nV_{n}$. In this paper, we attempt to find a natural definition of volume and perimeter in normed spaces $\mathbb{R}^{n}$ based precisely on the preservation of this Euclidean ratio $P(B)/V(B)=n$ for the unit ball $B\subset\mathbb{R}^{n}$. To achieve this, we introduce an intrinsic, recursive approach, starting with $n=2$ and using the definition of self-perimeter.
@@
The concept of self-perimeter is a central object of study in asymmetric Minkowski geometry and the theory of convex distance functions. A fundamental extremal problem in this field involves bounding the extremal self-perimeter of convex bodies in dimension two. The historical starting point is Golab’s theorem [6], which established that for centrally symmetric convex bodies, the self-perimeter is bounded between six and eight (with the maximum attained by a parallelogram, the minimum by the perfect hexagon, while the Euclidean $2\pi$ lies in between). Extending this problem to the space of asymmetric convex bodies expands the upper bound. The precise result states that the minimum self-perimeter of any planar convex body (calculated with respect to the optimal interior reference point) is bounded between six and nine. The lower bound of six is attained if and only if the body is an affinely regular hexagon. The absolute maximum of nine is attained if and only if the convex body is a triangle, a property formally proven by Grünbaum [7]. These results are further supported by the strict convexity of the self-perimeter with respect to the location of the interior reference point [8]. For regular $k$-gons, as well as for the family of planar convex bodies possessing $2\pi/k$ rotational symmetry, the self-perimeter exhibits periodic behavior depending on the remainder of $k$ modulo four [9, 10]. The literature demonstrates that when $k\equiv 0\pmod{4}$, the regular polygon constitutes a maximum (upper bound) for the family of shapes with the same degree of symmetry. Conversely, when $k\equiv 2\pmod{4}$, the regular polygon serves as a minimum [11]. This modular partition dictates the nature of the convergence of the self-perimeter to the limit of $2\pi$ (the value for a circle).
Extension to higher dimensions
@@
In this paper, we introduce a twist on Busemann’s definition [2] of an affine invariant perimeter of a convex set $K$ containing the origin in its interior:
@@
$\sigma_{K}^{(n-1)}(\nu_{x})\equiv\frac{\omega^{(n-1)}}{{\cal H}_{n-1}(K\cap\Sigma^{n-1}(\nu_{x}))}$ (2)
$\Sigma^{n-1}(\nu_{x})$ is the (n1)(n-1)-hyperplane perpendicular to the normal $\nu_{x}\in\partial K$ and
$\omega^{(n-1)}=\frac{\pi^{(n-1)/2}}{\Gamma\left(\frac{n-1}{2}+1\right)}$
is the volume of the unit (n1)(n-1)-ball in Euclidean space. Considering a centrally symmetric convex set $B\in\mathbb{R}^{n}$, we are looking for a sensible definition of a perimeter of a set $K\subset\mathbb{R}^{n}$ corresponding to the norm $\|\cdot\|_{B}$ induced by $B$ as a unit ball. Our definition is inspired by the affine invariant definition of Busemann (1) in dimension 22. In fact, if we consider $K$ to be the unit ball $B$ in $\mathbb{R}^{2}$, then equation (1) is the actual perimeter of $B$ in the metric $\|\cdot\|_{B}$ (see Sec. 3), called the self-perimeter of $B$.
@@
In our approach, we replace $\omega^{(n-1)}$ in $\sigma_{K}(\nu_{x})$ (2) by a recursive definition of the volume $\omega^{(n-1)}_{B}(\nu)$, depending on $B$ and on the unit vector $\nu\in\mathbb{R}^{n}$. See Definition 4.1 below.
Thus, the perimeter of a set $K\subset\mathbb{R}^{n}$ consistent with the norm $\|\cdot\|_{B}$ is:
$|\partial K|_{B}:=\int_{\partial K}\omega^{(n-1)}_{B}(\nu_{x})d{\cal H}_{n-1}(x)\$
while its affine invariant perimeter is an adaptation of (1, 2):
$P(K)=\int_{\partial K}\frac{\omega^{(n-1)}_{B}(\nu_{x})}{{\cal H}_{n-1}(K\cap\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)\ ,$
@@
This definition coincides with Busemann’s definition for $n=2$ since $\omega^{(1)}=2$ is independent of direction in $\mathbb{R}^{2}$. However, it replaces Busemann’s preservation of the Euclidean volume of the unit ball with the preservation of the Euclidean ratio $P(B)/\omega_{B}^{(n)}=n$ of perimeter/volume of the unit ball in $n$ dimensions (Definition 4.1).
@@
In 2D, the uniquely defined perimeter (measured via the norm) sets the area base case as $V_{2}=P_{2}/2$. For dimensions $n\geq 3$, the surface area is defined by integrating over the Euclidean sphere $S^{n-1}$, normalizing each surface element by the (n1)(n-1)-dimensional volume of the unit ball’s central cross-section orthogonal to the specified direction. It is remarkable that while Busemann’s definition (1) does not show any connection with the Minkowski norm, it does give the self-perimeter when applied to the unit ball of the same norm in dimension 2. This seems to be just a coincidence, due to the ansatz $V(B^{(1)})=\omega^{(1)}=2$.
@@
As we will show, the exact cancellation of the Jacobian in this formulation ensures that this definition of volume is strictly preserved under origin-preserving affine transformations. We extend the definition of $\omega_{B}^{(n)}$ to non-centrally symmetric convex bodies $B$, calculate explicitly the self-volume/perimeter of an $n$-simplex $\Delta^{n}$, and prove its convexity with respect to the chosen center. This implies the convexity of the self-perimeter/volume of any convex body in $\mathbb{R}^{n}$ composed of Cartesian products of simplices, 1D segments, and 2D convex sets. At the end of Section 4 we pose the following conjecture (see Section 4.4):
Conjecture
  • The CCS (convex centrally symmetric) body in $\mathbb{R}^{n}$ of maximal self-volume is the $n$-cube $\omega^{(n)}_{B}=2^{n}$.
  • The non-centrally symmetric, convex body in $\mathbb{R}^{n}$ of maximal self-volume with respect to its centroid is the $n$-simplex $B=\Delta^{n}$ ($\omega^{(n)}_{B}=\frac{(n+1)^{n}}{n!}$).
  • If $n=2k$, then the CCS body of minimal self-volume is the Cartesian product of $k$ hexagons ($\omega^{(n)}_{B}=3^{k}$). If $n=2k+1$, it is the Cartesian product of $k$ hexagons and a 1D interval ($\omega^{(n)}_{B}=2\times 3^{k}$).
@@
Beyond the foundational task of defining volume and surface area, this framework allows us to pose a novel inverse problem. In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively [4]. Within the context of our intrinsic definitions, we introduce an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set [1, 5]. Given that our perimeter and volume are defined recursively by integrating over the essential boundary—normalized by the self-volume of (n1)(n-1)-dimensional cross-sections—the natural inverse question arises: Given a positive measure $\mu$ on the Euclidean sphere $S^{n-1}$, does there exist a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\mu$?
In Section 5 we explore this problem in the two-dimensional case, providing perturbative solutions and identifying necessary conditions for its solvability around the Euclidean disc. In particular, we obtain that an $\mathcal{O}(\epsilon)$ perturbation of the uniform measure on the unit Euclidean circle yields an $\mathcal{O}(\epsilon^{1/2})$ perturbation of this circle, which depends only on the log 4-fold symmetric part of the measure (c.f., Theorem 5.1).
@@ 2 Fundamental definitions
Let $B\subset\mathbb{R}^{n}$ be a convex, centrally symmetric set (CCS). The norm $\|\cdot\|_{B}$ is defined in the usual way as:
@@
Let $K$ be a starlike set in $\mathbb{R}^{n}$ with respect to the origin. For $\theta\in S^{(n-1)}$ set:
$r_{K}(\theta)=\sup\{t>0;t\theta\in K\}\$
@@
If $K$ is centrally symmetric (CCS), then $r_{K}(\theta)=r_{K}(-\theta)$.
Definition $2.$
@@
3 Self-perimeter in the 2-plane
If $n=2$, there is a natural notion of a perimeter $P_{B}$ of $B$ as measured with respect to this norm. Indeed, we can use the norm $\|\cdot\|_{B}$ as a natural definition of length of an arc.
In the case $n=2$, we can introduce an explicit formulation for the perimeter:
@@ 3 Self-perimeter in the 2-plane
where dsds is the Euclidean length of the boundary of $B$, and $\vec{\tau}(s)$ is the tangent vector to the boundary at $s\in\partial B$. An explicit expression in terms of an integral on the Euclidean circle $\mathbb{S}^{1}$ is:
@@
The self-perimeter of a convex $B\ni\{0\}$ in $\mathbb{R}^{2}$ is:
$P(B)=\int_{0}^{2\pi}\frac{\sqrt{r^{2}_{B}(\phi)+(r^{{}^{\prime}}_{B}(\phi))^{2}}d\phi}{r_{B}\left(\phi+\arctan\left(\frac{r_{B}(\phi)}{r_{B}^{{}^{\prime}}(\phi)}\right)\right)}$ (4)
where $r_{B}(\phi):=r_{B}(\cos(\phi),\sin(\phi))$.
@@
The vector $\vec{r}=(r\cos(\phi),r\sin(\phi))$. Its tangent is $d\vec{r}/d\phi=(r^{{}^{\prime}}\cos(\phi)-r\sin(\phi),r^{{}^{\prime}}\sin(\phi)+r\cos(\phi))$. The slope of this direction is:
@@
The self-perimeter of CCS is always between 4 and 6.
Corollary 3.1.
If $r_{B}(\phi)$ is the radius function of a CCS in $\mathbb{R}^{2}$, then (4) takes values between 6 and 8.
Note that for $B\in CCS(\mathbb{R}^{2})$, $r_{B}$ is a $\pi$-periodic function.
The definition (4) can easily be extended to convex, non-centrally symmetric sets. Let $K$ be such a set and $p\in Int(K)$.
@@
If $B$ is convex but not centrally symmetric, the perimeter $P$ may depend on the orientation of the integral on the boundary. We agree on the positive (counter-clockwise) orientation for this definition.
Remark 3.2.
In addition, there is no upper limit to the self-perimeter for non-centrally symmetric sets. Indeed, if $p\in B$ is very close to $\partial B$, then evidently the self-perimeter of $B-\{p\}$ can be very large.
The following result was established in [8].
@@
Let $K\subset\mathbb{R}^{2}$ be a bounded convex set and $p\in Int(K)$. Then $p\rightarrow P(K-p)$ is strongly convex in $Int(K)$. In particular, there exists a unique interior point $p\in\text{int}(K)$ that minimizes the self-perimeter $P(K-\{p\})$.
For the sake of completeness, we provide a proof below. The self-perimeter of $K$ with respect to an interior point $p$ is defined by the integral:
$P(K-p)=\int_{\partial K}\frac{1}{r_{K,p}(\vec{\tau}(s))}ds$
@@
Hence, $r_{K,p}(v)$ is a concave function of pp.
@@
Since the self-perimeter $P(p)$ is an integral (a continuous sum) of strictly convex functions over the boundary $\partial K$ with respect to a positive length measure dsds, the function $P(p)$ itself is strictly convex on $\text{int}(K)$.